Write the equation of the line L satisfying the given geometric conditions.
L has y-intercept (0, -3) and is parallel to the line with equation y = 2/3x + 1.

Answer by Septimus314(7)

Equation of a line is y=mx + b
m = slope and b= y-intercept
here, b= -3, as that is your y-intercept. parallel lines have the same slope, so your slope is m=2/3, the same as in the line they gave.
This is all the information we need
y = 2/3x -3